OpenStudy (anonymous):

Calculus Limits at Infinity Please help? find lim h(x) as x approaches infinity, if possible. f(x)=  5x^2 - 3x + 7 a) h(x)= f(x)/x b) h(x)= f(x)/x^2 c) h(x)= f(x)/x^3

OpenStudy (anonymous):

a) make factors of given polynomial then divide by x, set limit and get answer

OpenStudy (abb0t):

You factor out the largest term. So it would be x\(^2\)

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

so is it like this? f(x)= x^2(5) (3) +7 ?

OpenStudy (abb0t):

\(\sf x^2[5-\frac{3}{x} +\frac{7}{x^2}]\)

OpenStudy (anonymous):

and do I have to put all that over x?

OpenStudy (cwrw238):

for b) h(x) = 5 - 3/x + 7/x^2

OpenStudy (cwrw238):

so what do you think the limit of 5 - 3/x + 7/x^2 is, as x approaches infinity?

OpenStudy (anonymous):

How do I solve that? am I supposed to take the square root of everything to cancel out?

OpenStudy (cwrw238):

no plug in x = infinity into h(x)

OpenStudy (cwrw238):

what does the value of 3/x approach as x approaches infinity?

OpenStudy (anonymous):

it approaches 3+ right?

OpenStudy (cwrw238):

no:- 3 / infinity is close to what value?

OpenStudy (cwrw238):

can you see that its zero?

OpenStudy (cwrw238):

the same goes for 7/x^2 so limit as x --> infinity of 5 - 3/x + 7/x^2 is 5

OpenStudy (anonymous):

I don't understand this!! :(

OpenStudy (anonymous):

so was I setting a) up correctly?

OpenStudy (cwrw238):

in the case of a) h(x) = f(x)/x so h(x) = 5x - 3 + 7/x now you consider the limit of this as x ---> infinity

OpenStudy (anonymous):

ok so I don't have to factor

OpenStudy (anonymous):

yeah I have to find the limit of 5x-3+7

OpenStudy (anonymous):

and that is 5x-3+7=0 5x+4= 0 x=4/5 ?

OpenStudy (kainui):

Turn the fraction into each of its individual terms, then find the limit of each term individually. So maybe you have (6x^2-x)/(2x^2) then make it: (6x^2/2x^2)+(-x/2x^2) and see that you have 3-(1/2x) and if you're taking the limit as that approaches infinity, the first term isn't dependent on x, so you have 3+ something else. The 1/(2x) as that approaches infinity approaches 0, so your final answer will be 3+0, or 3 for this example.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thnx